OpenStudy (anonymous):

What is the sum of a 11–term geometric series if the first term is 4 and the last term is 4,096? Answer a 9,124 b 8,188 c 11,264 d 45,056

OpenStudy (anonymous):

the geometric sequence is 2^n starting at n =2

OpenStudy (anonymous):

huh??

OpenStudy (anonymous):

so you have 2^2 = 4 2^3 = 8 2^4 = 16 2^5 = 32 2^6 = 64 it goes all the way until 2^12 get the sum of those 11 numbers